OpenStudy (anonymous):

What is the function rule for the table? x y 1 0.5 4 2 12 6 50 25

OpenStudy (anonymous):

what do you do to make each x = y

OpenStudy (anonymous):

no you divide but how much?/

OpenStudy (anonymous):

y=(?)x y equals x times what? .5= 1 times ??

OpenStudy (anonymous):

2=4 times ??

OpenStudy (anonymous):

y = 1. 5x

OpenStudy (anonymous):

you can also do x /2

OpenStudy (anonymous):

if you look to the table the corresponding values of x are the half of values of x. it means that y=kx here you have to guess such value of k that that kx become half of x.

OpenStudy (anonymous):

y equals one point five times x y equals one-half times x y equals x minus zero point five y equals x minus six Which one?

OpenStudy (anonymous):

y equals one-half times x

OpenStudy (anonymous):

like i told you y= kx if you put k= 1/2 then y= (1/2)x its the required relation.